# An Introduction to Decentralized Training and Execution in Cooperative Multi-Agent Reinforcement Learning

### v3

@@ 2 Decentralized, value-based methods
Value-based methods (such as Q-learning (Watkins and Dayan, 1992)) learn a value function and the policy is implicitly defined based on those values (i.e., choosing the action that maximizes the value function). Some of the earliest MARL methods are based on applying Q-learning to each agent. For instance, Claus and Boutilier (1998) decompose MARL approaches into independent learners (IL) and joint-action learners (JAL). ILs learn Q-functions that depend on their own information (ignoring other agents) while JALs learn joint Q-functions, which include the actions of the other agents. These concepts were originally defined for the fully observable, bandit (i.e., stateless) case but can be extended to the partially observable, stateful case. For example, we can define an IL Q-function as $Q_{i}({h}_{i},{a}_{i})$, which provides a value for agent $i$’s history ${h}_{i}$ and action ${a}_{i}$. A JAL Q-function $Q_{i}(\mathbf{h}{},\mathbf{a})$ for agent $i$ would depend on the joint information $\mathbf{h}{}$ and $\mathbf{a}$. All decentralized value-based MARL methods are ILs since they never have access to the joint information, $\mathbf{h}{}$ and $\mathbf{a}$.
In particular decentralized value-based MARL methods for Dec-POMDPs:
  • assume at each step, each agent, $i$, observes the action it took, ${a}_{i}$, its resulting observation, ${o}_{i}$, as well as the joint reward, $r$, and
  • estimate a value function based solely on this information (i.e., episodes consisting of sequences of ${a}_{i},{o}_{i},r$).
As a result, as we will discuss, any (partially observable) single-agent MARL method can be a DTE MARL method by ignoring the fact that other agents exist and learning as if they were the only agent in the problem. Such methods may have issues with convergence and performance (as we’ll discuss below) but this is exactly what is done by our first approach: Independent Q-learning.
2.1 Independent Q-learning (IQL)
@@
IQL just applies Q-learning to each agent independently. That is, each agent learns its own local Q-function, $Q_{i}$, using Q-learning on its own data. Pseudocode for a version of IQL for the Dec-POMDP case is provided in Algorithm 1. Just like Q-learning in the single-agent case, that algorithm uses a learning rate, $\alpha$, and some form of exploration, which in this case is $\epsilon$-greedy. Each agent initializes their Q-function (e.g., to 0, optimistically, etc.), and then iterates over episodes (either a fixed set or until convergence). During each episode, the agent chooses an action with exploration, sees the joint reward and its own observation and then updates the Q-value at the current history, ${h}_{i}$, given the current Q-estimates and the reward using the standard Q-learning update (line 9). The history is updated by appending the action taken and observation seen to the current history and the episode continues until the horizon is reached.
@@
Note that the Q-functions are inherently timed as the history length defines the current time step, making the algorithm fit with the finite-horizon case. Regardless, the algorithm can easily be extended to the episodic/indefinite-horizon case by including terminal states (or histories) instead of a fixed horizon.
1:  set $\alpha$ and $\epsilon$
@@
This algorithm runs in the underlying Dec-POMDP but the details of those dynamics in the environment are not shown (as they are not visible to the agent). For example, at the beginning of each episode, the state, $s$, is sampled from the initial state distribution ${{{b}}_{0}}$, and when agents take joint action $\mathbf{a}$ the reward is generated from the reward function, $R(s,\mathbf{a})$, the next state $s^{\prime}$ is sampled from the transition dynamics with probability $\Pr(s^{\prime}|s,\mathbf{a})$, and the joint observation is sampled from the observation function $\Pr(\mathbf{o}|s^{\prime},\mathbf{a})$. This process continues until the end of the episode.
While independent Q-learning ignores the other agents, it still depends on them in order to generate the joint reward $r$ and the local observation $o_{i}$ since both of these depend on the actions of all agents. That is, from agent $i$’s perspective, it is learning in a (single-agent) POMDP, as Algorithm 1 shows (the states of the POMDP would be the states of the Dec-POMDP plus the histories of the other agents).444 The idea of fixing other agent policies and then solving the resulting POMDP (either by planning methods (Nair et al., 2003) or RL methods (Banerjee et al., 2012)) has had success and results in a best response for agent $i$. Furthermore, iterating over agents, $i$, fixing other agent policies, $\neg i$, and calculating best responses will lead to a local optimum of the Dec-POMDP (Banerjee et al., 2012). Unfortunately, such methods require coordination to communicate which agent’s turn it is to learn so they are not appropriate for the DTE case.
In particular, as pointed out in previous papers (Lauer and Riedmiller, 2000) [^5] Note the paper considers the fully observable deterministic transition case but we extend the idea to the Dec-POMDP case., independent Q-learning is actually learning a Q-function for each agent based on the actions selected by the other agents during training. That is, agent $i$ would use data that depends on the actions taken by the other agents (and observations other agents see) even if it doesn’t directly observe those actions (and observations). In the partially observable case, the following Q-function would be learned:
$Q_{i}({h}_{i},{a}_{i})=\sum_{\mathbf{a}\in\mathbf{\mathbb{A}}}\hat{P}(\mathbf{a},\mathbf{h}{}|{h}_{i},{a}_{i})\left[r+\gamma\sum_{{o}_{i}}\hat{P}({o}_{i}|\mathbf{h}{},\mathbf{a})\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})\right]$ (1)
@@
In the case of IQL, the underlying POMDP that is being learned is non-stationary since the other agents are also learning and thus changing their policies over time. This may cause IQL to not converge (Tan, 1993). Convergence of Q-learning in the multi-agent case is an active area of research (Claus and Boutilier, 1998; Tuyls et al., 2003; Wunder et al., 2010; Hussain et al., 2023), it is still an open question what assumptions will allow convergence and to what (unlike the policy gradient case where convergence to a local optimum is assured under mild assumptions (Peshkin et al., 2000; Lyu et al., 2023)) [^6] In fact, while Q-learning can provably converge to a global optimum in the single-agent case (Jaakkola et al., 1993), there are no known MARL algorithms that have guaranteed convergence to a global optimum. This is an interesting area for future research!.
@@
2.2 Improving the performance of IQL
Several extensions of IQL have been developed to improve performance and better fit with the MARL setting. While the below algorithms were developed for the fully observable case (the multi-agent MDP (Oliehoek and Amato, 2016)), we extend them to the partially observable case here to fit better with our Dec-POMDP setting.
2.2.1 Distributed Q-learning
In order to promote coordination, Distributed Q-learning (Lauer and Riedmiller, 2000) makes strong optimistic assumptions. In particular, extending to the partially observable case, it uses the following Q-update:
$Q_{i}({h}_{i},{a}_{i})=\max\left\{Q_{i}({h}_{i},{a}_{i}),r+\gamma\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})\right\}$ (2)
where the current Q-value is kept if it is higher than (or equal to) the new estimate $r+\gamma\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})$. Otherwise, the Q-value is updated to be the new estimate. The intuition behind this optimistic update is that other agents will make mistakes during training because they have suboptimal policies or because they are exploring. Agents don’t necessarily want to learn from these mistakes. Instead, agents want to learn from the best policies of the other agents such as those that choose actions that generate high-reward values, even if they don’t happen very often. In the special case of deterministic MMDPs (the multi-agent fully observable case, see Oliehoek and Amato (2016)), this approach will learn the same Q-function as centralized Q-learning (which will converge to the optimal Q-values in the limit). This is unlike IQL which can get stuck in local optima or never converge at all.
@@ 2.2.1 Distributed Q-learning
In the more general stochastic case, distributed Q-learning is not able to distinguish between environmental stochasticity and agent stochasticity. That is, if getting a high-valued reward is unlikely (due to reward or transition stochasticity) distributed Q-learning will assume it can be received deterministically using max rather than taking the expectation (over other agent policies and stochastic transitions) as in Equation 1. As a result, distributed Q-learning is overly optimistic in the stochastic case and may perform poorly. Learning and incorporating the probabilities in Equation 1 and then calculating the expectation over the next step Q-value could fix this problem but would assume agents observe the actions of other agents which is not the case for decentralized training.
2.2.2 Hysteretic Q-learning
@@
$Q_{i}({h}_{i},{a}_{i})=\begin{cases}Q_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\\ Q_{i}({h}_{i},{a}_{i})\quad\text{else}\end{cases}$ (3)
@@
Hysteretic Q-learning, in contrast, adds the second learning rate $\beta$ to the negative experiences:
$Q_{i}({h}_{i},{a}_{i})=\begin{cases}Q_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\\ Q_{i}({h}_{i},{a}_{i})+\beta\delta\quad\text{else}\end{cases}$ (4)
@@ 2.2.3 Lenient Q-learning
${Q}_{i}({h}_{i},{a}_{i})=\begin{cases}{Q}_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\quad\text{or}\quad rand\sim U(0,1)>1-e^{-K*T({h}_{i},{a}_{i})}\\ {Q}_{i}({h}_{i},{a}_{i})\quad\quad\quad\text{else}\end{cases}$ (5)
@@
Lenient Q-learning can outperform hysteretic Q-learning since it can adjust the amount of optimism in a more fine-grained way but it also requires maintaining (and updating) temperature values as well as Q-values. Maintaining these values can be more problematic in the partially observable case since the history-action space grows exponentially with the problem horizon (or history length more generally).
2.3 Deep extensions, issues, and fixes
@@
In order to scale to larger domains, deep extensions of the above algorithms have been developed. These approaches are typically based on deep Q-networks (DQN) (Mnih et al., 2015). While DQN can scale to larger action and observation spaces, the decentralized multi-agent context can cause problems. We first discuss the basics of DQN and its extension to the partially observable case, DRQN, and then discuss a deep version of IQL, independent DRQN (Tampuu et al., 2017), the coordination issues with the approach, and some proposed fixes.
2.3.1 DQN and DRQN
Deep Q-networks (DQN) (Mnih et al., 2015) is an extension of Q-learning (Watkins and Dayan, 1992) to include a neural net as a function approximator. That is, DQN learns $Q_{\theta}(s,a)$, parameterized with $\theta$ (i.e., $\theta$ represents the parameters of the neural network), by minimizing the loss:
$\mathcal{L}(\theta)=\mathbb{E}_{<s,a,r,s^{\prime}>\sim\mathcal{D}}\Big{[}\big{(}y-Q_{\theta}(s,a)\big{)}^{2}\Big{]}\text{, where\,\, }y=r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(s^{\prime},a^{\prime})$ (6)
which is just the squared TD error—the difference between the current estimated value, $Q_{\theta}(s,a)$, and the new value gotten from adding the newly seen reward to the previous Q-estimate at the next state, $Q_{\theta^{-}}(s^{\prime},a^{\prime})$. Because learning neural networks can be unstable, a separate target action-value function, $Q_{\theta^{-}}$, and an experience replay buffer $\mathcal{D}$ (Lin, 1992) are implemented to stabilize learning. The target network is an older version of the Q-estimator that is updated periodically, the experience replay buffer stores $s,a,s^{\prime},r$ sequences and single $s,a,s^{\prime},r$ tuples are i.i.d. sampled for updates. As shown in Figure 2(a), the neural network (NN) outputs values for all actions ($a\in\mathbb{A}$) to make maxing at a state possible with a single forward pass (rather than iterating through the actions).
Deep recurrent Q-networks (DRQN) (Hausknecht and Stone, 2015) extends DQN to handle partial observability, where some number of recurrent layers (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) are included to maintain an internal hidden state which is an abstraction of the history (as shown in Figure 2(b)). Because the problem is partially observable, $o,a,o^{\prime},r$ sequences are stored in the replay buffer during execution. The update equation is very similar to that of DQN:
$\mathcal{L}(\theta)=\mathbb{E}_{<h,a,r,o>\sim\mathcal{D}}\Big{[}\big{(}y-Q_{\theta}(h,a)\big{)}^{2}\Big{]}\text{, where\,\, }y=r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(h^{\prime},a^{\prime})$ (7)
but since a recurrent neural network (RNN) is used, the internal state of the RNN can be thought of as a history representation. That is, instead of learning Q-functions using full histories (i.e., $h=o_{0},a_{1},o_{1},r_{1},\ldots,o_{t-1},a_{t},o_{t},r_{t}$), $h$ is the internal state of the RNN after sequentially inputing the history sequence ($h=o_{0},a_{1},o_{1},\ldots,o_{t-1},a_{t},o_{t}$). Similarly, $h^{\prime}=hao^{\prime}$, which can be thought of as the new history after appending $o^{\prime}$ and $a$ to $h$, which is achieved by inputing $a$ and $o^{\prime}$ to the RNN when its current state is $h$. Technically, a whole history (e.g., episode) should be sampled from the replay buffer to train the recurrent network but it is common to use a fixed history length (e.g., sample $o,a,o^{\prime},r$ sequences of length 10). As mentioned above, many implementations just use observation histories rather than full action-observation histories. With the popularity of DRQN, it has become common to add recurrent layers to standard (fully observable) deep reinforcement learning methods when solving partially observable problems.
(a) DQN
(b) DRQN
(c) Independent DRQN
Figure 2: DQN, DRQN, and independent DRQN diagrams.
2.3.2 Independent DRQN (IDRQN)
@@
Independent DRQN (IDRQN) (Tampuu et al., 2017) [^7] Again, the original algorithm only considers the fully observable case but we consider the extension to the Dec-POMDP case. combines IQL and DRQN. As seen in Algorithm 2, the basic structure is the same as IQL but ${Q}_{i}$ is estimated using an RNN parameterized by $\theta$ (shown in Figure 2(c)), along with using a target network, $\theta^{-}$, and a replay buffer, $\mathcal{D}$. First, episodes are generated by interacting with the environment (e.g., $\epsilon$-greedy exploration), which are added to the replay buffer. This buffer typically has a fixed size and new data replaces old data when the buffer is full. Episodes can then be sampled from the replay buffer (either single episodes or as a minibatch of episodes). In order to have the correct internal state, RNNs are trained sequentially. Training is done from the beginning of the episode until the end [^8] It is common in practice to use a fixed history length and sample windows of that length rather than training with full episodes., calculating the internal state and the loss at each step and updating $\theta$ using gradient descent. The target network $\theta^{-}$ is updated periodically (every $C$ episodes in the code) by copying the parameters from $\theta$.
1:  set $\alpha$, $\epsilon$, and $C$
@@
6:     ${h}_{i}\leftarrow\theta^{RNN}$initial history is the initial RNN state of the network}
@@
15:        ${h}_{i}\leftarrow\theta^{RNN}$initial history is the initial RNN state of the network}
@@
While the pseudocode uses a horizon, the resulting Q-functions no longer reflect the time step based on the history since histories of different lengths may be combined into the same representation in the RNN. As a result, untimed/stationary Q-values are learned that just depend on the RNN internal state, which we still call $h$, and not the time step. This stationary representation is technically incorrect for the finite-horizon case (even though it is widely used in practice). Again, the code can be adjusted to the episodic/indefinite-horizon case by replacing the horizon with terminal states or the infinite-horizon case by removing the loop over episodes (both of which fix the issue with timed Q-functions as these cases do not require them). Lastly, the pseudocode, for this approach and throughout the text, is written to be simple and highlight the main algorithmic components. There are many implementation tricks, especially in the deep case, that are not included.
@@
Independent DRQN can perform well but has a number of issues. For example, early methods disabled experience replay (Foerster et al., 2016) [^9] Other methods for improving experience replay in MARL exist, but they typically assume the CTDE setting (such as Foerster et al. (2017)).. They found that performance was worse with experience replay than without it. This is because agents are no longer learning concurrently when using the replay buffer—experience is generated from the behavior of all agents (i.e., concurrently) but is then put into the buffer for learning. When learning, different agents will sample and learn from different experiences. Learning from this different data makes performance noisy and unstable. This makes intuitive sense since agents will update their Q-functions (and resulting policies) based on (potentially very) different experiences. We discuss a simple fix next.
2.3.3 Deep hysteretic-Q and concurrent buffer
Hysteretic Q-learning has been extended to the deep case in the form of decentralized hysteretic deep recurrent Q-networks (Dec-HDRQN) (Omidshafiei et al., 2017). Dec-HDRQN makes two main contributions: the use of hysteresis with D(R)QN, and a concurrent replay buffer to improve performance [^10] The paper also discusses the multi-task case, but we don’t discuss that aspect here..
@@ 2.3.3 Deep hysteretic-Q and concurrent buffer
The concurrent buffer is also relatively straightforward. Concurrent experience replay trajectories (CERTs), are based on the idea of creating a joint experience replay buffer and then sampling from that buffer. In particular, we would like to have a joint replay buffer with joint experience tuples, ${\mathbf{a}}^{e}_{1},{\mathbf{o}}^{e}_{1},{r}^{e}_{1},\ldots,{\mathbf{a}}^{e}_{t},{\mathbf{o}}^{e}_{t},{r}^{e}_{t}$, where the superscript here represents the episode number, $e$. Agents could then sample from this joint buffer to get the local information from the same episode and time step: ${h}_{i},{a}_{i},{o}_{i},r$ (where ${h}_{i}$ comes from the internal state of agent $i$’s RNN after passing the local history of actions and observations until the time step being sampled). Of course, in a decentralized learning context, agents will not have access to such a joint buffer, but agents can store their part of the joint buffer by indexing by episode $e$ and timestep $t$. That is, as seen in Figure 3, agents store their local trajectories along axes for $e$ and $t$. When an $e$ and $t$ is sampled, each agent $i$ will sample data from the same episode and timestep, producing the same data as would be produced with the joint buffer. As long as agents have the same size buffers and coordinate on the seeds for the random number generator, this approach can be used for decentralized learning.
Figure 3: Concurrent experience replay trajectories (CERTs) (from (Omidshafiei et al., 2017))
2.3.4 Deep lenient Q-learning

### v4

@@ 3.2 Decentralized, value-based methods
As discussed in Section 1.3.1, value-based methods (such as Q-learning (Watkins and Dayan, 1992)) learn a value function and the policy is implicitly defined based on those values (i.e., choosing the action that maximizes the value function). Some of the earliest MARL methods are based on applying Q-learning to each agent. For instance, Claus and Boutilier (1998) decompose MARL approaches into independent learners (IL) and joint-action learners (JAL). ILs learn Q-functions that depend on their own information (ignoring other agents) while JALs learn joint Q-functions, which include the actions of the other agents. These concepts were originally defined for the fully observable, bandit (i.e., stateless) case but can be extended to the partially observable, stateful case. For example, we can define an IL Q-function as ${Q}_{i}({h}_{i},{a}_{i})$, which provides a value for agent $i$’s history ${h}_{i}$ and action ${a}_{i}$. A JAL Q-function ${Q}_{i}(\mathbf{h}{},\mathbf{a})$ for agent $i$ would depend on the joint information $\mathbf{h}{}$ and $\mathbf{a}$. All decentralized value-based MARL methods are ILs since they never have access to the joint information, $\mathbf{h}{}$ and $\mathbf{a}$ and many of the algorithms use the ‘independent’ term in their name.
I first discuss the tabular (Q-learning based) methods, starting with a simple application of single-agent Q-learning and then discuss improvements to learning for the multi-agent setting. Next, I discuss extensions to the deep case (DQN-based), again in terms of a simple application of DRQN as well as issues and fixes for the multi-agent case.
3.2.1 Independent Q-learning (IQL)
@@
IQL just applies Q-learning to each agent independently, treating other agents as part of the environment. That is, each agent learns its own local Q-function, $Q_{i}$, using Q-learning on its own data. Pseudocode for a version of IQL for the Dec-POMDP case is provided in Algorithm 1. Just like Q-learning in the single-agent case, that algorithm uses a learning rate, $\alpha$, and some form of exploration, which in this case is $\epsilon$-greedy. Each agent initializes their Q-function (e.g., to 0, optimistically, etc.), and then iterates over episodes (either a fixed set or until convergence). During each episode, the agent chooses an action with exploration, the algorithm sees the joint reward and the agent’s own observation and then the algorithm updates the Q-value at the current history, ${h}_{i}$, given the current Q-estimates and the reward using the standard Q-learning update (line 9). The history is updated by appending the action taken and observation seen to the current history and the episode continues until the horizon is reached.
@@
The Q-functions are inherently tied to the history length, which defines the current time step, making the algorithm suitable for finite-horizon cases. Regardless, the algorithm can easily be extended to the episodic/indefinite-horizon case by including terminal states (or histories) instead of a fixed horizon.
1:  set $\alpha$ and $\epsilon$ (learning rate, exploration)
@@
This algorithm runs in the underlying Dec-POMDP but the details of those dynamics in the environment are not shown (as they are not visible to the agent and algorithm). For example, at the beginning of each episode, the state, $s$, is sampled from the initial state distribution ${{{b}}_{0}}$, and when agents take joint action $\mathbf{a}$ the reward is generated from the reward function, $R(s,\mathbf{a})$, the next state $s^{\prime}$ is sampled from the transition dynamics with probability $\Pr(s^{\prime}|s,\mathbf{a})$, and the joint observation is sampled from the observation function $\Pr(\mathbf{o}|s^{\prime},\mathbf{a})$. This process continues until the end of the episode.
While independent Q-learning ignores the other agents, it still depends on them in order to generate the joint reward $r$ and the local observation $o_{i}$ since both of these depend on the actions of all agents. That is, from agent $i$’s perspective, it is learning in a (single-agent) POMDP, as Algorithm 1 shows (the states of the POMDP would be the states of the Dec-POMDP plus the histories of the other agents).111 The idea of fixing other agent policies and then solving the resulting POMDP (either by planning methods (Nair et al., 2003) or RL methods (Banerjee et al., 2012)) has had success and results in a best response for agent $i$. Furthermore, iterating over agents, $i$, fixing other agent policies, $\neg i$, and calculating best responses will lead to a local optimum of the Dec-POMDP (Banerjee et al., 2012). Unfortunately, such methods require coordination to communicate which agent’s turn it is to learn so they are not appropriate for the DTE case.
In particular, as pointed out in previous papers (Lauer and Riedmiller, 2000) [^2] Note the paper considers the fully observable deterministic transition case but I extend the idea to the Dec-POMDP case., independent Q-learning is actually learning a Q-function for each agent based on the actions selected by the other agents during training. That is, agent $i$ would use data that depends on the actions taken by the other agents (and observations other agents see) even if it doesn’t directly observe those actions (and observations). In the partially observable case, the following Q-function would be learned:
$Q_{i}({h}_{i},{a}_{i})=\sum_{\mathbf{a}\in\mathbf{\mathbb{A}}}\hat{P}(\mathbf{a},\mathbf{h}{}|{h}_{i},{a}_{i})\left[r+\gamma\sum_{{o}_{i}}\hat{P}({o}_{i}|\mathbf{h}{},\mathbf{a})\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})\right]$ (3.1)
@@
In the case of IQL, the underlying POMDP that is being learned is nonstationary since the other agents are also learning and thus changing their policies over time. This may cause IQL to not converge (Tan, 1993). Convergence of Q-learning in the multi-agent case is an active area of research (Claus and Boutilier, 1998; Tuyls et al., 2003; Wunder et al., 2010; Hussain et al., 2023), it is still an open question what assumptions will allow convergence and to what (unlike the policy gradient case where convergence to a local optimum is assured under mild assumptions (Peshkin et al., 2000; Lyu et al., 2023)) [^3] In fact, while Q-learning can provably converge to a global optimum in the single-agent case (Jaakkola et al., 1993), there are no known MARL algorithms that have guaranteed convergence to a global optimum. This is an interesting area for future research!.
@@
3.2.2 Improving the performance of IQL
Several extensions of IQL have been developed to improve performance and better fit with the MARL setting. While the below algorithms were developed for the fully observable case (the multi-agent MDP described in Section 2, I extend them to the partially observable case here to fit better with our Dec-POMDP setting.
Distributed Q-learning
In order to promote coordination, Distributed Q-learning (Lauer and Riedmiller, 2000) makes strong optimistic assumptions, prioritizing high-reward outcomes from agent actions while discounting suboptimal ones during learning. In particular, extending to the partially observable case, it uses the following Q-update:
$Q_{i}({h}_{i},{a}_{i})=\max\left\{Q_{i}({h}_{i},{a}_{i}),r+\gamma\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})\right\}$ (3.2)
where the current Q-value is kept if it is higher than (or equal to) the new estimate $r+\gamma\max_{{a}_{i}^{\prime}}Q_{i}({h}_{i}^{\prime},{a}_{i}^{\prime})$. Otherwise, the Q-value is updated to be the new estimate. The intuition behind this optimistic update is that other agents will make mistakes during training because they have suboptimal policies or because they are exploring. Agents don’t necessarily want to learn from these mistakes. Instead, agents want to learn from the best policies of the other agents such as those that choose actions that generate high-reward values, even if they don’t happen very often. In the special case of deterministic MMDPs, this approach will learn the same Q-function as centralized Q-learning (which will converge to the optimal Q-values in the limit). This is unlike IQL which can get stuck in local optima or never converge at all.
@@ 3.2.2 Improving the performance of IQL
In the more general stochastic case, distributed Q-learning is not able to distinguish between environmental stochasticity and agent stochasticity. That is, if getting a high-valued reward is unlikely (due to reward or transition stochasticity) distributed Q-learning will assume it can be received deterministically using max rather than taking the expectation (over other agent policies and stochastic transitions) as in Equation 3.1. As a result, distributed Q-learning is overly optimistic in the stochastic case and may perform poorly. Learning and incorporating the probabilities in Equation 3.1 and then calculating the expectation over the next step Q-value could fix this problem but would assume agents observe the actions of other agents which is not the case for decentralized training.
Hysteretic Q-learning
@@
$Q_{i}({h}_{i},{a}_{i})=\begin{cases}Q_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\\ Q_{i}({h}_{i},{a}_{i})\quad\text{else}\end{cases}$ (3.3)
@@
Hysteretic Q-learning, in contrast, adds the second learning rate $\beta$ that is applied to the negative experiences:
$Q_{i}({h}_{i},{a}_{i})=\begin{cases}Q_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\\ Q_{i}({h}_{i},{a}_{i})+\beta\delta\quad\text{else}\end{cases}$ (3.4)
@@
${Q}_{i}({h}_{i},{a}_{i})=\begin{cases}{Q}_{i}({h}_{i},{a}_{i})+\alpha\delta\quad\text{if}\quad\delta>0\quad\text{or}\quad rand\sim U(0,1)>1-e^{-K*T({h}_{i},{a}_{i})}\\ {Q}_{i}({h}_{i},{a}_{i})\quad\quad\quad\text{else}\end{cases}$ (3.5)
@@
Lenient Q-learning can outperform hysteretic Q-learning by fine-tuning optimism levels on a per-state-action basis. Unfortunately, it also requires maintaining (and updating) temperature values as well as Q-values. Maintaining these values can be more problematic in the partially observable case since the history-action space grows exponentially with the problem horizon (or history length more generally).
3.2.3 Deep extensions, issues, and fixes
@@
In order to scale to larger domains, deep extensions of the above algorithms have been developed. These approaches are typically based on deep Q-networks (DQN) (Mnih et al., 2015). While DQN can scale to larger action and observation spaces, the decentralized multi-agent context can cause challenges such as non-stationarity and coordination difficulties. I discuss a deep version of IQL, independent DRQN (Tampuu et al., 2017), the coordination issues with the approach, and some proposed fixes. Details about DQN and DRQN are in Section 1.3.1.
Figure 3.1: Independent DRQN diagram.
Independent DRQN (IDRQN)
@@
Independent DRQN (IDRQN) (Tampuu et al., 2017) [^4] Again, the original algorithm only considers the fully observable case but I consider the extension to the Dec-POMDP case. combines IQL and DRQN. As seen in Algorithm 2, the basic structure is the same as IQL but ${Q}_{i}$ is estimated using an RNN parameterized by $\theta$ (shown in Figure 3.1), along with using a target network, $\theta^{-}$, and a replay buffer, $\mathcal{D}$. First, episodes are generated by interacting with the environment (e.g., $\epsilon$-greedy exploration), which are added to the replay buffer. This buffer typically has a fixed size and new data replaces old data when the buffer is full. Episodes can then be sampled from the replay buffer (either single episodes or as a minibatch of episodes). In order to have the correct internal state, RNNs are trained sequentially. Training is done from the beginning of the episode until the end [^5] It is common in practice to use a fixed history length and sample windows of that length rather than training with full episodes., calculating the internal state and the loss at each step and updating $\theta$ using gradient descent. The target network $\theta^{-}$ is updated periodically (every $C$ episodes in the code) by copying the parameters from $\theta$.
1:  set $\alpha$, $\epsilon$, and $C$ (learning rate, exploration, and target update frequency)
@@
6:     ${h}_{i}\leftarrow\emptyset$initial history is empty}
@@
15:        ${h}_{i}\leftarrow\emptyset$
@@
Again, the code can be adjusted to the episodic/indefinite-horizon case by replacing the horizon with terminal states or the infinite-horizon case by removing the loop over episodes. The pseudocode, for this approach and throughout the text, is written to be simple and highlight the main algorithmic components. There are many implementation tricks, especially in the deep case, that are not included.
@@
Independent DRQN can perform well but has a number of issues. For example, early methods disabled experience replay (Foerster et al., 2016) [^6] Other methods for improving experience replay in MARL exist, but they typically assume the CTDE setting (such as Foerster et al. (2017)).. They found that performance was worse with experience replay than without it. This is because agents are no longer learning concurrently when using the replay buffer—experience is generated from the behavior of all agents (i.e., concurrently) but is then put into the agent’s buffer for learning. When learning, different agents will sample from their buffers separately, learning from different experiences. Learning from this different data makes performance noisy and unstable. This makes intuitive sense since agents will update their Q-functions (and resulting policies) based on (potentially very) different experiences. I discuss a simple fix next.
Deep hysteretic-Q and concurrent buffer
Hysteretic Q-learning has been extended to the deep case in the form of decentralized hysteretic deep recurrent Q-networks (Dec-HDRQN) (Omidshafiei et al., 2017). Dec-HDRQN makes two main contributions: the use of hysteresis with D(R)QN, and a concurrent replay buffer to improve performance. [^7] The paper also discusses the multi-task case, but I don’t discuss that aspect here.
@@
The concurrent buffer is also relatively straightforward. Concurrent experience replay trajectories (CERTs), are based on the idea of creating a joint experience replay buffer and then sampling from that buffer. In particular, we would like to have a joint replay buffer with joint experience tuples, ${\mathbf{a}}^{e}_{1},{\mathbf{o}}^{e}_{1},{r}^{e}_{1},\ldots,{\mathbf{a}}^{e}_{t},{\mathbf{o}}^{e}_{t},{r}^{e}_{t}$, where the superscript here represents the episode number, $e$. Agents could then sample from this joint buffer to get the local information from the same episode and time step: ${h}_{i},{a}_{i},{o}_{i},r$ (where ${h}_{i}$ represents the history at the current time step such as the internal state of agent $i$’s RNN after passing the local history of actions and observations until the time step being sampled). Of course, in a decentralized learning context, agents will not have access to such a joint buffer, but agents can store their part of the joint buffer by indexing by episode $e$ and timestep $t$. That is, as seen in Figure 3.2, agents store their local trajectories along axes for $e$ and $t$. When an $e$ and $t$ is sampled, each agent $i$ will sample data from the same episode and timestep, producing the same data as would be produced with the joint buffer. As long as agents have the same size buffers and coordinate on the seeds for the random number generator, this approach can be used for decentralized learning.
Figure 3.2: Concurrent experience replay trajectories (CERTs) (from (Omidshafiei et al., 2017))
Deep lenient Q-learning